### v1

# https://sarahweiii.github.io/meshlrm/ MeshLRM: Large Reconstruction Model for High-Quality Meshes

Abstract
We propose MeshLRM, a novel LRM-based approach that can reconstruct a high-quality mesh from merely four input images in less than one second. Different from previous large reconstruction models (LRMs) that focus on NeRF-based reconstruction, MeshLRM incorporates differentiable mesh extraction and rendering within the LRM framework. This allows for end-to-end mesh reconstruction by fine-tuning a pre-trained NeRF LRM with mesh rendering. Moreover, we improve the LRM architecture by simplifying several complex designs in previous LRMs. MeshLRM’s NeRF initialization is sequentially trained with low- and high-resolution images; this new LRM training strategy enables significantly faster convergence and thereby leads to better quality with less compute. Our approach achieves state-of-the-art mesh reconstruction from sparse-view inputs and also allows for many downstream applications, including text-to-3D and single-image-to-3D generation.
Keywords:
Sparse-view reconstruction High-quality mesh Large Reconstruction Models![[Uncaptioned image]](extracted/2404.12385v1/figures/teaser.png)
1 Introduction
High-quality 3D mesh models are the core of 3D vision and graphics applications and 3D editing, rendering, and simulation tools are specifically optimized for them. 3D mesh assets are usually created manually by expert artists or alternatively reconstructed from multi-view 2D images. Traditionally this has been done with complex photogrammetry systems [55, 19, 49, 50], though recent neural approaches, such as NeRF [41], offer a simpler end-to-end pipeline through per-scene optimization. However, these neural methods typically produce volumetric representations [41, 42, 6, 33, 78, 63]; converting them into meshes requires additional optimization in post [59, 77, 70, 47]. Furthermore, both traditional and neural approaches require a large number of input images and long processing time (up to hours), limiting their applicability in time-sensitive design workflows.
Our goal is efficient and accurate 3D asset creation via few-shot mesh reconstruction with direct feed-forward network inference and no per-scene optimization. We base our approach on recent large reconstruction models (LRMs) [23, 30, 73, 64] for 3D reconstruction and generation. Existing LRMs use triplane NeRFs [5] as the 3D representation for high rendering quality. While these NeRFs can be converted into meshes via a Marching Cube (MC) [39] post-processing, this leads to a significant drop in rendering quality and geometric accuracy.
We address this with MeshLRM, a novel transformer-based large reconstruction model, designed to directly output high-fidelity 3D meshes from sparse-view inputs. Specifically, MeshLRM incorporates differentiable surface extraction and rendering into a NeRF-based LRM. We apply a recent Differentiable Marching Cube (DiffMC) [70] technique to extract an iso-surface from the triplane NeRF’s density field and render the extracted mesh using a differentiable rasterizer [29]. This enables end-to-end training of MeshLRM using mesh rendering losses, to optimize it to produce high-quality meshes that render realistic images.
We train MeshLRM by initializing the model using volumetric NeRF rendering; since our meshing components do not introduce any new parameters they can start from these pre-trained weights. However, we find that training a mesh-based LRM with differentiable MC is still highly challenging. The primary issue lies in the (spatially) sparse gradients from the DiffMC operation, that only affect surface voxels and leave the vast empty space untouched. This leads to poor local minima for model optimization and manifests as floaters in the reconstructions. We address this with a novel ray opacity loss that ensures that empty space along all pixel rays maintains near-zero density, effectively stabilizing the training and guiding the model to learn accurate floater-free surface geometry. Our end-to-end training approach reconstructs high-quality meshes with rendering quality that surpasses NeRF volumetric rendering (see Tab. 5).
Our DiffMC-based meshing technique is general and can potentially be applied to any NeRF-based LRMs for mesh reconstruction. In this work, we present a simple and efficient LRM architecture comprising of a large transformer model that directly processes concatenated multi-view image tokens and triplane tokens with purely self-attention layers to regress final triplane features for NeRF and mesh reconstruction. In particular, we simplify many complex design choices used in previous LRMs [30, 73, 64], including the removal of pre-trained DINO modules in image tokenization and replacing the large triplane decoder MLP with small two-layer ones. We train MeshLRM on the Objaverse dataset [14] with a novel low-resolution pre-training and high-resolution fine-tuning strategy. These design choices lead to a state-of-the-art LRM model with faster training and inference and higher reconstruction quality.
In summary, our key contributions are:
A novel LRM-based framework that integrates differentiable mesh extraction and rendering for end-to-end few-shot mesh reconstruction.
A novel ray opacity loss for stabilizing DiffMC-based generalizable training.
An efficient LRM architecture and training strategies to enable fast and high-quality reconstruction.
2 Related Work
Mesh Reconstruction. Despite the existence of various 3D representations, meshes remain the most widely used format in industrial 3D engines. Reconstructing high-quality meshes from multi-view images is a long-standing challenge in computer vision and graphics. Classically, this is addressed via a complex multi-stage photogrammetry pipeline, integrating techniques such as structure from motion (SfM) [55, 1, 49], multi-view stereo (MVS) [19, 50], and mesh surface extraction [39, 27]. Recently, deep learning-based methods have also been proposed to address these problems for higher efficiency and accuracy [62, 58, 75, 24, 25, 11]. However, the photogrammety pipeline is not only costly, requiring dense input images and long processing time, but also often suffers from low-quality mesh rendering. Our approach enables sparse-view mesh reconstruction through direct feed-forward network inference in an end-to-end manner. We leverage neural volumetric representations and differentiable rendering, training our model to output high-quality meshes that can render realistic images.
Neural Reconstruction. Neural rendering techniques have recently gained significant attention for their ability to produce high-quality 3D reconstructions for realistic rendering [80, 37, 41, 28]. Most recent methods are based on NeRF [41] and reconstruct scenes as various volumetric radiance fields [41, 6, 18, 42, 72] with per-scene rendering optimization. Though radiance fields can be converted into meshes with Marching Cubes [39], the quality of the resulting mesh is not guaranteed. Previous methods have aimed to transform radiance fields into implicit surface representations (SDFs) [63, 76, 43], enhancing mesh quality from Marching Cubes. In addition, some methods [77, 70, 59, 47] attempt to directly extract meshes from radiance fields using differentiable surface rendering. However, these methods all require dense input images and long per-scene optimization time, which our approach avoids.
On the other hand, generalizable neural reconstruction has been explored, often leveraging MVS-based cost volumes [7, 38, 26] or directly aggregating multi-view features along pixels’ projective epipolar lines [79, 65, 57]. While enabling fast and few-shot reconstruction, these methods can only handle small baselines, still requiring dense images with local reconstruction in overlapping windows to model a complete object. More recently, transformer-based large reconstruction models (LRMs) [23, 30, 64, 73] have been proposed and achieved 3D NeRF reconstruction from highly sparse views. In this work, we propose a novel efficient LRM, incorporating DiffMC techniques [70], to enable high-quality sparse-view mesh reconstruction via direct feed-forward inference.
3D Generation. Generative models have seen rapid progress with GANs [21] and, more recently, Diffusion Models [56] for image and video generation. In the context of 3D generation, many approaches utilize 3D or multi-view data to train 3D generative models with GANs [71, 22, 20, 5] or diffusion models [2, 73, 67, 8, 36], which are promising but limited by the scale and diversity of 3D data. DreamFusion [45] proposed to leverage the gradient of a pre-trained 2D diffusion model to optimize a NeRF for text-to-3D generation with a score distillation sampling (SDS) loss, achieving diverse generation results beyond common 3D data distributions. This approach inspired many follow-up approaches, to make the optimization faster or improve the results quality [31, 68, 60, 10, 40]. However, these SDS-based methods rely on NeRF-like per-scene optimization, which is highly time-consuming, often taking hours. More recently, a few attempts have been made to achieve fast 3D generation by utilizing pre-trained 2D diffusion models to generate multi-view images and then perform 3D reconstruction [30, 34, 35]. In particular, Instant3D [30] leverages an LRM to reconstruct a triplane NeRF from four sparse generated images. In this work, we focus on the task of sparse-view reconstruction and propose to improve and re-target a NeRF LRM to directly generate meshes with higher quality. Our model can be naturally coupled with a multi-view generator, such as the ones in [30, 53], to enable high-quality text-to-3D and single-image-to-3D generation (see Fig. 5 and Fig. 6).
3 Method
In this section, we present our MeshLRM that reconstructs high-quality meshes within 1 second. We start with describing our backbone transformer architecture (Sec. 3.1), which simplifies and improves previous LRMs. We then introduce our two-stage framework for training the model: we first (Sec. 3.2) train the model to predict NeRF from sparse input images by supervising volume renderings at novel views, followed by refining mesh surface extractions (Sec. 3.3) by performing differentiable marching cubes on the predicted density field and minimizing a surface rendering loss with differentiable rasterization.
3.1 Model Architecture
As shown in Fig. 2, we propose a simple transformer-based architecture for MeshLRM, mainly consisting of a sequence of self-attention-based transformer blocks over concatenated image tokens and triplane tokens, similar to PF-LRM [64]. In contrast to PF-LRM and other LRMs [23, 30, 73], we simplify the designs for both image tokenization and triplane NeRF decoding, leading to fast training and inference. We next describe the model in detail.
Input posed image tokenization. MeshLRM adopts a simple tokenizer for posed images, inspired by ViT[16]. We convert the camera parameters for each image into Plücker ray coordinates [44] and concatenate them with the RGB pixels (3-channel) to form a 9-channel feature map. Then we split the feature map into non-overlapping patches, and linearly transform them as transformer’s input. With this process, the model does not need additional positional embedding as in ViT, since Plücker coordinates contain spatial information.
Note that our image tokenizer is much simpler than previous LRMs that use a pre-trained DINO ViT [4] for image encoding. We empirically find that the pre-trained DINO ViT is unnecessary, possibly because DINO is mainly trained for intra-view semantic reasoning, while 3D reconstruction mainly requires inter-view low-level correspondences. By dropping this per-view DINO encoding, the model also enables a better flow between raw pixels and 3D-related information.
Transformer. We concatenate multi-view image tokens and learnable triplane (positional) embeddings, and feed them into a sequence of transformer blocks [61], where each block is comprised of self-attention and MLP layers. We add layer normalization [3] before both layers (i.e., Pre-LN architecture), and use residual connections. This deep transformer network enables comprehensive information exchange among all the tokens, and effectively models intra-view, inter-view, and cross-modal relationships. The output triplane tokens, contextualized by all input views, are then decoded into the renderable triplane NeRF. The output image tokens are dropped. Lastly, each triplane token is unprojected with a linear layer and further unpatchified to triplane features via reshaping. 111This operator is logically the same to a shared 2D-deconvolution with stride and kernel on each plane. All the predicted triplane features are then assembled into the final triplane NeRF.
Tiny density and color MLPs. Previous LRMs [23, 30, 73, 64] all use a heavy shared MLP (e.g., 9 hidden layers with a hidden dimension of 64) to decode densities and colors from triplane features. This design leads to slow rendering in training. This is a bottleneck for MeshLRM since we need to compute a dense density grid, using the MLP and triplane features, for extracting the mesh with DiffMC. Therefore, we opt to use tiny MLPs with a narrower hidden dimension of and fewer layers. In particular, we use an MLP with one hidden layer for density decoding and another MLP with two hidden layers for color decoding. Compared to a large MLP, our tiny MLPs can lead to speed-up in training without compromising quality (see Tab. 4). We use separate MLPs since the density MLP and color MLP are used separately in the Marching Cubes and surface rendering processes, respectively. We empirically find this MLP separation is necessary for DiffMC fine-tuning (described later in Sec. 3.3) as it largely improves the optimization stability which is critical to large-scale training.
While being largely simplified, our transformer model can effectively transform input posed images into a triplane NeRF for density and color decoding. The density and color are then used to achieve both radiance field rendering for 1st-stage volume initialization (Sec. 3.2) and surface extraction+rendering for 2nd-stage mesh reconstruction (Sec. 3.3).
3.2 Stage 1: Efficient Training for Volume Rendering
We train our model with ray marching-based radiance field rendering (as in NeRF [41]) to bootstrap our model, providing good initialized weights for mesh reconstruction. Instead of training directly using high-res () input images like prior LRM works [23, 30], we develop an efficient training scheme, inspired by ViT [16]. Specifically, we first pretrain our model with images until convergence and then finetune it for fewer iterations with images. This training scheme cuts the computing cost significantly, hence obtaining much better quality than training at 512-res from scratch given the same amount of compute (see Tab. 1).
256-res pretraining. The pre-training uses 256-resolution images for both model input and output. We use a batch size of 8 objects per GPU and sample 128 points per ray during ray marching. In contrast to training with 512-res images from scratch (like previous LRMs), our training efficiency benefits from the low-res pre-training from two factors: shorter sequence length for computing self-attention and fewer samples per ray for volume rendering (compared to our high-res fine-tuning).
512-res finetuning. For high-res fine-tuning, we use 512-resolution images for the model input and output. We use a batch size of 2 per GPU and densely sample 512 points per ray, which is the number in low-res pretraining. Here, we compensate for the increased computational costs (from longer token sequences and denser ray samples) by reducing the batch size 4 times, achieving a similar training speed (per iteration) to the low-res pretraining. It’s also worth noting that our resolution-varying pretraining and finetuning scheme benefits from the use of Plücker coordinates for camera conditioning. It avoids the positional embedding interpolation [16, 46] and has inherent spatial smoothness.
Loss. We use an L2 regression loss and a perceptual loss (proposed in [9]) to supervise the renderings from both phases. Since rendering full-res images is not affordable for volume rendering, we randomly sample a patch from each target 256- or 512-res image for supervision with both losses. We also randomly sample 4096 pixels per target image for additional L2 supervision, allowing the model to capture global information beyond a single patch. The loss for volume rendering training is expressed by:
where we use for both 256- and 512-res training.
3.3 Stage 2: Stable Training for Surface Rendering
Once trained with volume rendering, our model already achieves high-fidelity sparse-view NeRF reconstruction, which can be rendered with ray marching to create realistic images. However, directly extracting a mesh from the NeRF’s density field results in significant artifacts as reflected by Tab. 2. Therefore, we propose to fine-tune the network using differentiable marching cubes and differentiable rasterization, enabling high-quality feed-forward mesh reconstruction.
Mesh extraction and rendering. We compute a density grid by decoding the triplane features and adopt a recent differentiable marching cubes (DiffMC) technique [70] to extract mesh surface from the grid. The DiffMC module is based on a highly optimized CUDA implementation, much faster than existing alternatives [52, 51], enabling fast training and inference for mesh reconstruction.
To compute the rendering loss we need to render our generated mesh. We do so using a differentiable rasterizer Nvdiffrast [29] and utilize the triplane features to (neurally) render novel images from our extracted meshes. This full rendering process is akin to deferred shading where we first obtain per-pixel XYZ locations via differentiable rasterization before querying the corresponding triplane features and regressing per-pixel colors using the color MLP. We supervise novel-view renderings with ground-truth images, optimizing the model for high-quality end-to-end mesh reconstruction.
However, using the rendering loss alone leads to high instability during training and severe floaters in the meshes (see Fig. 3). This is mainly caused by the sparsity of density gradients during mesh rendering: unlike volume rendering which samples points throughout the entire view frustum, mesh (rasterization) rendering is restricted to surface points, lacking gradients in the empty scene space beyond the surface.
Ray opacity loss. To stabilize the training and prevent the formation of floaters, we propose to use a ray opacity loss. This loss is applied to each rendered pixel ray, expressed by:
where represents the ground truth surface point along the pixel ray, is randomly sampled along the ray between and camera origin, and is the volume density at ; when no surface exists for a pixel, we sample inside the object bounding box along the ray and use the far ray-box intersection as .
In essence, this loss is designed to enforce the empty space in each view frustum to contain near-zero density. Here we minimize the opacity value , computed using the ray distance from the sampled point to the surface. This density-to-opacity conversion functions as an effective mechanism for weighting the density supervision along the ray with lower loss values for points sampled closer to the surface. An alternative approach would be to directly regularize the density using an L1 loss, but we found this to lead to holes in surfaces as all points contribute the same, regardless of their distance to the surface.
Combined losses. To measure the visual difference between our renderings and ground-truth (GT) images, we use an L2 loss and a perceptual loss proposed in [9] similar to stage 1. To compute the ray opacity loss , we obtain surface points using the GT depth maps. In addition, to further improve our geometry accuracy and smoothness, we apply an L2 normal loss to supervise the face normals of our extracted mesh with GT normal maps in foreground regions. Our final loss for mesh reconstruction is
where we use , and in our experiments. Since mesh rasterization is significantly cheaper than volume ray marching, we render the images at full resolution (e.g., in our experiment) for supervision, instead of the random patches+rays used in Stage 1 volume training.
4 Experiments
4.1 Dataset and Evaluation Protocols
Our model is trained on the Objaverse dataset [14] (730K objects) for the 1st-stage volume rendering training and is subsequently fine-tuned on the Objaverse-LVIS subset (46K objects) for the 2nd-stage surface rendering finetuning. Empirically, the Objaverse-LVIS subset has higher quality and previous work [48, 69, 13] shows that fine-tuning favors quality more than quantity. We thus follow this common practice. We evaluate the reconstruction quality of MeshLRM alongside other existing methods on the GSO [17], NeRF-Synthetic [41], and OpenIllumination [32] datasets, employing PSNR, SSIM, and LPIPS as metrics for rendering quality and bi-directional Chamfer distance (CD) as the metric for mesh geometry quality. For the CD metric, since we cannot expect the model to precisely reconstruct unseen parts, we cast rays from all test views and sample 100K points at the ray-surface intersections for each object. The Chamfer distance is computed using the points sampled from both the ground-truth and reconstructed mesh within a bounding box, measured in unit. All results are generated using four sparse-view input images.
4.2 Analysis and Ablation Study
Volume rendering (Stage 1) training strategies. To verify the effectiveness of our training strategy that uses 256-res pretraining and 512-res fine-tuning (details in Sec. 3.2), we compare with a model having the same architecture but trained with high-resolution only (i.e., 512-res from scratch). Tab. 1 shows the quantitative results on the GSO dataset with detailed training settings and timings used for the two training strategies. As seen in the table, with the same total compute budget of 64 hours and 128 GPUs, our low-to-high-res training strategy achieves significantly better performance, with a 2.6dB increase in PSNR. The key to enabling this is our fast low-res pretraining, which takes only 2.6 seconds per iteration, allowing for many more iterations to be trained within a shorter period and enabling much faster convergence. Benefiting from effective pretraining, our high-res finetuning can be trained with a smaller batch size that reduces the iteration time. In general, our 1st-stage training strategy significantly accelerates the LRM training, leading to high-quality NeRF reconstruction. This naturally improves the mesh reconstruction quality in the second stage.
PSNR | SSIM | LPIPS | #Pts/Ray | BatchSize | #Iter | |||
512-res from scratch | 25.53 | 0.892 | 0.123 | 512 | 1024 | 7.2s | 32k | 64hrs |
256-res pretrain | 28.13 | 0.923 | 0.093 | 128 | 1024 | 2.6s | 60k | 44hrs |
512-res fine-tune | 512 | 256 | 3.6s | 20k | 20hrs |
Effectiveness of surface fine-tuning (Stage 2). We justify the effectiveness and significance of our 2nd-stage surface fine-tuning by comparing our final meshes with those directly extracted from 1st-stage model using marching cubes on GSO dataset. The quantitative results are shown in Tab. 2, and the qualitative results are presented in the Appendix. Note that, while our 1st-stage NeRF reconstruction, i.e. MeshLRM (NeRF), can achieve high volume rendering quality, directly extracting meshes from it with Marching Cubes (MC), i.e. ‘MeshLRM (NeRF)+MC’, leads to a significant drop of rendering quality on all metrics. On the other hand, our final MeshLRM model, fine-tuned with DiffMC-based mesh rendering, achieves significantly better mesh rendering quality and geometry quality than the MeshLRM (NeRF)+MC baseline, notably increasing PSNR by 2.5dB and lowering CD by 0.58. This mesh rendering quality is even comparable — with a slight decrease in PSNR but improvements in SSIM and LPIPS — to our 1st-stage volume rendering results of MeshLRM (NeRF).
PSNR | SSIM | LPIPS | CD | |
MeshLRM (NeRF) | 28.13 | 0.923 | 0.093 | - |
MeshLRM (NeRF)+MC | 25.48 | 0.903 | 0.097 | 3.26 |
MeshLRM | 27.93 | 0.925 | 0.081 | 2.68 |
PSNR | SSIM | LPIPS | CD | |
w/o Ray Opacity Loss | 18.46 | 0.836 | 0.218 | 51.98 |
w/o Normal Loss | 27.93 | 0.925 | 0.080 | 2.87 |
Ours (full model) | 27.93 | 0.925 | 0.081 | 2.68 |
Surface fine-tuning losses. We demonstrate the significance of the different losses applied during the mesh fine-tuning stage (details in Sec. 3.3) with an ablation study in Tab. 3 and Fig. 3. We observe a significant performance drop in the model without our proposed ray opacity loss, leading to severe floater artifacts as shown in Fig. 3. This is because, without supervision on empty space, the model may produce floaters to overfit to the training views rather than learning the correct geometry, as it does not receive a penalty for inaccuracies in empty space. The normal loss particularly helps to generate better geometry, i.e. a lower Chamfer distance. While the normal loss does not lead to quantitative improvements in rendering (metrics are about the same) on the GSO dataset, we still observe qualitative improvements with this loss, especially when using generated images, as shown in Fig. 3 (right). Empirically, the normal loss leads to better robustness in handling inconsistent input views, a common challenge in text-to-3D or image-to-3D scenarios (e.g. the cactus). Moreover, the geometric improvements from the normal loss are also crucial for real applications to use our meshes in 3D engines (like the one in Fig. 1), where accurate shapes are critical for high-fidelity physically-based rendering.
PSNR | SSIM | LPIPS | ||
Triplane+Big MLP | 27.44 | 0.915 | 0.081 | 3.6s |
Triplane+Tiny MLP | 27.47 | 0.915 | 0.080 | 2.7s |
Tiny MLPs. To justify our choices of using tiny MLPs for triplane coding, instead of a large one in previous LRMs, we compare with a version that replaces the two tiny MLPs with a 10-layer (i.e., 9 hidden layers) 64-width shared MLP decoder (as used in [23, 30]). The quantitative results on the GSO dataset are in Tab. 4. We find that the tiny MLP can achieve similar performance to the large MLP while offering a notable training speed advantage (2.7s / step vs 3.6s / step). Moreover, we observe that the heavy shared MLP hardly converge during the Stage 2 DiffMC training while the tiny separate MLPs could.
4.3 Comparison Against Baselines
Render | FPS | Params | Quality (GSO) | ||||||
(GPUday) | PSNR | SSIM | LPIPS | CD | |||||
In3D-LRM [30] | 128 pts/ray | 0.5 | 0.07s | 500M | 1287 | 26.54 | 0.893 | 0.064 | - |
MeshLRM (NeRF) | 512 pts/ray | 2 | 0.06s | 300M | 1282.7 | 28.13 | 0.923 | 0.093 | - |
In3D-LRM [30]+MC | Mesh | >60 | 1s | 500M | 1287 | 23.70 | 0.875 | 0.111 | 3.40 |
MeshLRM | >60 | 0.8s | 300M | 1283 | 27.93 | 0.925 | 0.081 | 2.68 |
Inference | OpenIllumination | NeRF-Synthetic | ||||||
Time | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | CD | |
FreeNeRF [74] | 3hrs | 12.21 | 0.797 | 0.358 | 18.81 | 0.808 | 0.188 | - |
ZeroRF [54] | 25min | 24.42 | 0.930 | 0.098 | 21.94 | 0.856 | 0.139 | 6.01 |
MeshLRM | 0.8s | 26.10 | 0.940 | 0.070 | 21.85 | 0.850 | 0.137 | 4.94 |
Comparisons with feed-forward methods. To the best of our knowledge, the LRM model proposed in Instant3D [30] is the only previous model that can also achieve 3D object reconstruction from four sparse views. As reported by Instant3D, previous cost-volume-based feed-forward methods (like SparseNeus[38]) cannot handle this challenging case. Therefore, we focus on comparing with the LRM model in Instant3D (labeled as In3D-LRM) on the GSO dataset, showing quantitative results with detailed training settings and timings in Tab. 5. In the table, we show our results from both stages and compare them with both In3D-LRM and In3D-LRM + MC; for In3D-LRM + MC, Marching Cubes is directly applied to extract meshes from their NeRF reconstruction.
As shown in Tab. 5, our model, across both stages, achieves superior quality compared to Instant3D. As expected, In3D-LRM + MC leads to much worse rendering quality than In3D-LRM, due to the quality drop caused by MC. Thanks to our effective 2nd-Stage DiffMC-based training, MeshLRM can reconstruct significantly better meshes in terms of both rendering quality (with 4.2dB PSNR and 0.05 SSIM improvements) and geometry quality (lowering the CD by 0.72). Our mesh rendering is even better than their volume rendering (In3D-LRM), notably increasing their PSNR by 1.4dB and SSIM by 0.03, while achieving substantially faster rendering speed because of outputting meshes. Our superior quality is also reflected by the qualitative comparisons in Fig. 4, where our mesh renderings reproduce more texture and geometric details than the baselines. From Tab. 5, we can also see that our high reconstruction quality is even achieved with a smaller model size and substantially less compute, less than half of the total GPUxDay compute cost required for Instant3D-LRM. This is enabled by our simplified LRM architecture designs and efficient low-res-to-high-res 1st-stage training strategy, leading to significantly faster training speed and convergence. Besides, our model also leads to a faster inference speed, enabling high-quality mesh reconstruction within 1 second. Overall, our MeshLRM leads to state-of-the-art sparse-view mesh reconstruction quality and achieves high efficiency in parameter utilization, and training, inference, and rendering speed. Meanwhile, our MeshLRM (NeRF) has good performance as well and is a state-of-the-art LRM model for sparse-view NeRF reconstruction.
Comparisons with per-scene optimization methods. We also compare our MeshLRM with recent per-scene optimization methods that are designed for sparse-view NeRF reconstruction, including ZeroRF [54] and FreeNeRF [74]. Because of using per-scene optimization, these methods are much slower, requiring tens of minutes and even several hours to reconstruct one single scene. As a result, it is not practical to evaluate them on the GSO dataset, comprising more than 1000 objects. Therefore, we conduct this experiment on the NeRF-Synthetic and OpenIllumination datasets, following the settings used in ZeroRF [54] with four input views. Tab. 6 shows the quantitative novel view synthesis results, comparing our mesh rendering quality with their volume rendering quality. Since ZeroRF also offers a post-processing step to improve their mesh reconstruction from Marching Cubes, we also compare our mesh geometry with theirs using the CD metric on the NeRF-Synthetic dataset (where GT meshes are accessible).
As shown in Tab. 6, our NeRF-Synthetic results outperform FreeNeRF significantly; we also achieve comparable rendering quality and superior geometry quality to ZeroRF. On the challenging OpenIllumination real dataset, our approach outperforms both FreeNeRF and ZeroRF by a large margin, consistently for all three metrics. The OpenIllumination dataset, consisting of real captured images, demonstrates our model’s capability to generalize to real captures for practical 3D reconstruction applications, despite being trained only on rendered images. In addition to our superior (or comparable) quality, our feed-forward mesh reconstruction approach is significantly faster than these optimization-based NeRF methods, in terms of both reconstruction (1000 to 10000 speed up) and rendering.
5 Applications
Our approach achieves efficient and high-quality mesh reconstruction in less than one second, thus naturally enabling many 3D generation applications, when combined with multi-view image generation techniques.
Text-to-3D Generation We apply the multi-view diffusion models from Instant3D [30] to generate 4-view images from text inputs, followed by our Mesh-LRM to achieve text-to-3D generation. Our results and comparison with the original Instant3D pipeline with In3D-LRM (that is already quantitatively compared in Tab. 5 in terms of a reconstruction task) are shown in Fig. 5. Our approach leads to better mesh quality and fewer rendering artifacts.
Image-to-3D Generation We apply Zero123++ [53] v1.2 to generate multi-view images from a single image input, followed by our MeshLRM to achieve image-to-3D generation. Our results and comparison with the One2345++ [34], which uses a diffusion-based reconstruction model, are shown in Fig. 6. Our results are of much higher quality, containing sharper textures and more faithful geometric details than One2345++.
6 Conclusion
In this paper, we have presented MeshLRM, a novel Large Reconstruction Model that can directly output high-quality meshes. We achieved it by applying the Differentiable Marching Cubes (DiffMC) method and differentiable rasterization to fine-tune a pre-trained NeRF-based LRM trained with volume rendering. As DiffMC requires backbone efficiency, we brought multiple improvements (tiny shared MLP and simplified image tokenization) to the LRM architecture, facilitating both NeRF and mesh training. We also found that a low-to-high-res training strategy can significantly accelerate the training of the NeRF-based model. Compared with existing methods, our method has both quality increase and speed improvement and is the only one that can output high-quality meshes. In addition, we showed that our method can be directly applied to applications like text-to-3D and image-to-3D generation. As meshes are the most widely accepted format for 3D assets in the industry, we believe our method takes a step towards automating 3D asset creation and potentially opens up new possibilities with novel 3D workflows.
Acknowledgement
Thanks to Nathan Carr for his support on this project and to Linghao Chen for helping to run the One2345++ baseline.
Appendix 0.A Effectiveness of surface fine-tuning (Stage 2).
We have discussed the effectiveness of our second-stage fine-tuning in Sec. 4.2 with quantitative results shown in Tab. 2 in the main paper. We now show examples of qualitative results in Fig. 7, comparing our final meshes (MeshLRM) with the meshes (MeshLRM (NeRF) + MC) generated by directly applying Marching Cubes on the 1st-stage model’s NeRF results. As shown in the figure, the MeshLRM (NeRF) + MC baseline can lead to severe artifacts with non-smooth surfaces and even holes, whereas our final model with the surface rendering fine-tuning can effectively address these issues and produce high-quality meshes and realistic renderings. These qualitative results demonstrate the large visual improvements achieved by our final model for mesh reconstruction, reflecting the big quantitative improvements shown in the paper.
Appendix 0.B ABO and OpenIllumination datasets
ABO | OpenIllumination | |||||
PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | |
In3D-LRM [30] | 27.50 | 0.896 | 0.093 | 8.96 | 0.568 | 0.682 |
MeshLRM (NeRF) | 28.31 | 0.906 | 0.108 | 20.53 | 0.772 | 0.290 |
In3D-LRM [30]+MC | 22.11 | 0.850 | 0.144 | 8.92 | 0.507 | 0.691 |
MeshLRM | 26.09 | 0.898 | 0.102 | 20.51 | 0.786 | 0.218 |
We compare our proposed method with In3D-LRM [30] on 1000 randomly sampled examples in the ABO [12] dataset with quantitative results shown in Tab. 7. In particular, ABO is a highly challenging synthetic dataset that contains many objects with complex glossy materials. This is a big challenge for our model, as well as In3D-LRM’s model, since both assume Lambertian appearance. In this case, the NeRF models are often better than the mesh models, since NeRF rendering can possibly fake glossy effects by using colors from multiple points along each pixel ray while mesh rendering cannot. Nonetheless, as shown in the table, both our NeRF and final mesh model can outperform In3D-LRM’s NeRF and mesh variants respectively. Especially, thanks to our effective MC-based finetuning, our mesh rendering quality surpasses that of In3D-LRM + MC by a large margin (e.g. 3.92dB in PSNR and 0.48 in SSIM).
On the other hand, we also compare with In3D-LRM on the full OpenIllumination [32] dataset in Tab. 7. To clarify, this experiment setting is different from the setting used in Tab. 6 in the main paper, where we follow ZeroRF [54] to test only 8 objects with combined masks in OpenIllumination; here, we evaluate all models on the full dataset with 100 objects with object masks and we also take square crops that tightly bounding the objects from the rendered/GT images to compute rendering metrics, avoiding large background regions. As shown in the table, we observe that In3D-LRM cannot generalize to handle this challenging real dataset, leading to very low PSNRs, despite being trained on the same training data as ours. In contrast, our model can still work well on this out-of-domain dataset with high rendering quality.
Appendix 0.C Additional Implementation Details
The transformer size follows the transformer-large config [15]. It has layers and a model width of . Each attention layer has a attention head and each head has a dim of . The intermediate dimension of the MLP layer is . We use GeLU activation inside the MLP and use Pre-LN architecture. The layer normalization is also applied after the input image tokenization, and also before the triplane’s unpachifying operator.
For the training of both stages, we use AdamW with . Other Adam-specific parameters are set as default and we empirically found that the model is not sensitive to them. A weight decay of is applied to all parameters excluding the bias terms and layer normalization layers. The learning rate for stage 1 is . We also use cosine learning rate decay and a linear warm-up in the first 2K steps. For stage 2, we use a learning rate of , combined with cosine learning rate decay and a linear warm-up during the first 500 steps. In total, there are 10k fine-tuning steps. The resolution of DiffMC is within a bounding box, which is consistent with the triplane resolution.
Appendix 0.D Limitations
Since our model employs surface rendering and assumes only Lambertian appearance without performing any inverse rendering, it is not sufficiently robust when the input images contain complex materials, e.g. some examples in the ABO dataset. The generated mesh may use white color to fit the specular areas. We believe that incorporating inverse rendering into the current pipeline could address this issue. However, this requires sufficient training data with accurate material annotations, which we leave for future exploration. On the other hand, while handling highly sparse input views, our model requires input camera poses. Although poses can be readily obtained from text- or image-to-multi-view models [30, 53] for 3D generation tasks, calibrating sparse-view poses for real captures can be a big challenge. In the future, it will be an interesting direction to explore combining our approach with recent pose estimation techniques [64, 66] for joint mesh reconstruction and camera calibration.